# Online Transfer Learning for RSV Case Detection

### 03 Feb 2024 v1

@@ IV Dataset
The research data were sourced from the University of Pittsburgh’s Neptune Research Data Warehouse [40], and the study protocol received approval from the University of Pittsburgh Institutional Review Board (Protocol STUDY20030193). The dataset consists of 2,263,360 encounters from five EDs at the UPMC over nine years from January 2011 to May 2020. In regions with temperate climates, RSV infections predominantly manifest during the colder winter months. The onset of these infections is usually observed around October and November in the Northern Hemisphere. Typically, the highest incidence of RSV cases, or the peak period, occurs in January and February. As the season progresses, the frequency of RSV infections starts to decline, with the final cases of the season generally being recorded between March and May [11]. Given the seasonal nature of RSV outbreaks, we defined one season as spanning from June of one year to May of the following year, rather than using a calendar year division for data segmentation. Table I details the sizes of encounters for each season. An encounter was classified as related to RSV if it included one of any of the following International Classification of Diseases (ICD) codes at discharge:
@@
Encounters that did not include any of these ICD codes were categorized as non-RSV.
Our dataset contains 7614 features, including demographic information, lab test results, and natural language processing extracted clinical findings, detailed in Table II. The clinical findings were obtained by using the MetaMap parser [41], which applies pattern-matching and deduction rules. The clinical notes for each encounter serve as input for the parser, which extracts a set of clinical findings and their values. The values for each finding are either present (P), negated (N), or missing (M), where missing indicates that clinicians did not mention the finding in the note, according to the parser. In cases where an encounter had multiple clinical notes, we used the union of all clinical findings across all the notes. Contradictions across notes were resolved in favor of present findings, and this also applied to multiple occurrences of a finding in a single note.
@@
Figure 2: Change in weights over time.
Note that the x-axis is scaled non-uniformly to enhance data visibility.
Figure 2: Change in weights over time.
@@ V-A Experiment Settings
In this study, we employed an online multi-source transfer learning ensemble model to enhance our model’s adaptability in a specific target domain. This model uniquely incorporated datasets from previous seasons, each treated as a distinct source domain. Specifically, for targeting a particular season, such as 2019-2020, the model was trained on data from individual seasons preceding 2019-2020, with each season (2011 to 2018) considered separately. This approach ensures that each season’s unique data characteristics are captured effectively.
This approach mirrors real-world situations, where models must adjust to new and varied data from the target domain while leveraging knowledge from historical data. This is particularly useful in biomedical fields where labeled data may be limited, thereby improving the model’s ability to generalize and maintain robustness.
To more closely replicate real-world conditions, our model was set in an online learning environment for the target domain, updating continually with each new data point. This mimics the ongoing influx of medical data in actual scenarios, ensuring our model remains current and effective.
@@
We evaluated our model’s performance using the Area Under the Receiver Operating Characteristic Curve (AUROC), an important measure for gauging a model’s discriminative performance when classifying positive and negative cases across diverse data sequences. This offers a comprehensive assessment of the model’s overall effectiveness.
@@
In the specified algorithm, two hyperparameters, $\alpha$ and $\beta$, play crucial roles in the dynamic adjustment of model weights during the learning process. The setting of $\alpha$ and $\beta$ is consistent across all domains. The hyperparameter $\alpha$ is set to $\log_{2}10$ = 3.32, a value carefully selected after experimenting with several hyperparameter settings. This medium value for $\alpha$ was chosen to balance the weight penalty applied to a classifier’s weight upon making an incorrect prediction. This particular value for $\alpha$ strikes a balance between being too lenient and too punitive, allowing for a gradual decrease in a classifier’s influence while still retaining its contributions where they prove useful.
@@
The hyperparameter $\beta$, set at $1/200000$, plays a critical role in the volume-adaptive component of our algorithm. This parameter is instrumental in fine-tuning the adjustment of the target weight $w_{T}$, aligning it with the growing volume of incoming data. Here, the factor $j$ serves as an index representing the sequence of data points being processed. The specific choice of $1/200000$ for $\beta$ strikes a delicate balance. When the number of instances has not yet reached 200,000, this configuration ensures that the target weight is subtly suppressed, preventing premature dominance in the model’s decision-making process. However, once the count surpasses this threshold, the target weight is progressively amplified, reflecting the increased reliability and relevance of the more substantial data corpus. This mechanism ensures that the target model’s influence is appropriately scaled, growing more pronounced as the volume of target data becomes more significant, thereby sensitively responding to the increase in data volume while simultaneously tempering the rate of weight adjustment.
@@
By adopting this value for $\beta$, the algorithm ensures a balanced and responsive adaptation to the accumulating target data. Such a setup allows our ensemble model to systematically transition its focus, increasingly prioritizing the most recent data, which is presumed to be the most relevant for the current learning context. This approach exemplifies the nuanced and progressive nature of online learning, where the model dynamically adapts to the evolving data landscape, ensuring that its predictions remain accurate and pertinent.
@@ V-B Results and Analysis
Our primary objective is to develop a model adept at classifying data from the current season, which is data from the season 2019-2020 in this study. To achieve this, we trained individual models for each season ranging from 2011 to 2018. Subsequently, we evaluated the performance of these models using data from the season 2019-2020 to ascertain their effectiveness in handling present-season data. The results of these assessments are comprehensively tabulated and presented in Table III, providing a clear overview of each model’s performance across this temporal span.
@@
For the results, a clear upward trend in AUROC scores is evident as we approach the season of 2019-2020. This trend is especially pronounced in models trained with more recent data, indicating an improvement in the classifiers’ ability to distinguish between positive and negative cases as they are trained on data closer to the target season. This improvement can be attributed to the increasing relevance and recency of the data. As the training data gets closer to the target season, it likely becomes more representative of the current patterns and trends, thereby enhancing the classifiers’ predictive accuracy.
@@
To evaluate the performance of PPAW and PVAW strategies, we compare them with baseline methods, including:
  • Online Learning Model: This model employs an online learning strategy, focusing exclusively on sequentially acquired data from the target domain for training.
@@
  • Model Trained from All Previous Data: This approach involves pre-training a model using a combined dataset from all source domains.
@@
  • Ensemble Model with Equal Weights: This method employs an ensemble model that integrates various individual source models, assigning them equal and fixed weights throughout the process.
@@
  • Ensemble Model with Volume-based Weights: This strategy uses an ensemble model that weights individual source models according to the volume of their respective training datasets.
@@
  • Ensemble Model with Time-based Weights: This ensemble method integrates individual models based on their temporal proximity to the testing season (target domain), defining ‘time distance’ as the number of seasons separating a model’s training data from the testing period.
The final three approaches within the ensemble model framework employ predetermined weights, which are determined in one of three ways: evenly distributed across all models, adjusted according to the sizes of their training datasets, or allocated baspleed on the models’ temporal proximity to the target domain’s data period.
@@
In our comparative analysis of various online learning strategies against the backdrop of sequential epidemiological data, distinct performance metrics emerge, highlighting the efficacy and adaptability of these models as shown in Table IV. The Online Learning Model, with an AUROC of 0.850 (P$=$0.037), showcases a commendable capacity for leveraging sequentially acquired data, underscoring the viability of online learning in dynamically evolving domains. Slightly surpassing this, the Model Trained from All Previous Data, at an AUROC of 0.853 (P$<$0.001), demonstrates the potential advantages of a comprehensive, cross-domain dataset for pre-training, affirming the value of extensive historical data. Conversely, the Ensemble with Equal Weights method, despite its statistical significance, registers a lower AUROC of 0.728 (P$<$0.001), suggesting that uniform weighting across diverse models may dilute the predictive precision. More notably, the Ensemble with Volume-based Weights, at an AUROC of 0.680 (P$<$0.001), indicates that volume-centric model weighting may not effectively capture the intricate data nuances necessary for accurate target domain predictions. However, an improvement is observed with the Ensemble with Time-based Weights, achieving an AUROC of 0.761 (P$<$0.001), which implies that temporal relevance plays a crucial role in enhancing model performance, albeit not to the extent of the more sophisticated methods. Remarkably, both the Predictive Performance-Adaptive Weighting (PPAW) and the Predictive Volume-Adaptive Weighting (PVAW) strategies exhibit superior performance metrics, with PPAW matching the Model Trained from All Previous Data at an AUROC of 0.853 (P$<$0.001), and PVAW leading the cohort with an AUROC of 0.870. This delineates the PVAW strategy’s predominant advantage in dynamically calibrating model weights based on the evolving data landscape, marking a significant stride towards optimizing predictive models for healthcare applications, particularly in the nuanced domain of Respiratory Syncytial Virus (RSV) data analysis.
@@
TABLE IV: Comparative results of different models, featuring P values from DeLong’s test for statistical differences between their respective AUROCs and the AUROC of the PVAW method.
@@
Trained from All Previous Data 0.852 (P$<$0.001)
@@
PPAW 0.853 (P$<$0.001)
PVAW 0.870
Figure 3: Possibility of features present given RSV.
Figure 4: $Log_{10}LR$ of features present.
Our proposed PVAW approach, which utilizes dynamically updated weights, surpasses these conventional baseline methods. The essence of its enhanced performance lies in its ability to dynamically adapt, customizing its relevance to the current data. This dynamic adaptation is achieved by continuously adjusting the weights of the models based on their immediate performance with new data. Such a strategy enables the model to respond effectively to changing data trends and to identify patterns that were not apparent in earlier training datasets. As a result, our method shows a marked improvement in both accuracy and effectiveness in classifying medical data, outstripping the other methodologies evaluated in the study.
Fig. 2 illustrates the dynamic changes in the model weights corresponding to data from different seasons as new instances are introduced over time, providing a visual correlation with the tabulated performance metrics. Each line on the graph represents the weight trajectory for a model trained on data from a specific season, spanning from 2011 to 2018. Initially, all model weights start with a high value, denoting an equal influence within the ensemble. As the number of instances increases, the graph shows significant fluctuations in these weights. Notably, models corresponding to 2016-2017 and 2017-2018 maintain higher weights throughout, signifying their sustained predictive relevance over time. This trend aligns with Table III’s results, where the AUROC scores for 2016-2017 and 2017-2018 are the highest among the evaluated seasons. These scores indicate that the classifiers trained on 2016-2017 and 2017-2018 data are most effective at accurately predicting outcomes for the target season of 2019-2020. The congruence between the graph’s weight trends and the AUROC scores in Table III emphasizes the ensemble model’s capability to adaptively prioritize the most relevant historical data for current predictions, enhancing the overall accuracy of the RSV case detection model.
Fig. 3 and Fig. 4 present the possibility of features present in the RSV positive cases and the $Log_{10}$ likelihood ratio (LR) of variates respectively. 13 variates with the highest LR (with possibility$>$0.01) in our target season 2019-2020 were selected as the comparative features. In Fig. 3, the possibility of Age - less than 5 appears high with values ranging from 0.68 to 0.96 in all years, while other features only have around or below 0.2 for possibility. As shown in Fig. 4, with relatively small ranges of values, Age - less than 5 and Increased work of breathing are two much more steady features than others observed through all years. Compared to season 2019, the data points in seasons 2012-2013 and 2013-2014 have been the farthest from the 2019-2020 season, suggesting that the feature patterns in the 2012-2013 and 2013-2014 seasons are the least similar to our target season in 2019-2020, which also aligns with our model results in Table III that seasons 2012-2013 and 2013-2014 present the lowest AUROC. In contrast, features from seasons 2017-2018 and 2018-2019 are mainly located closest to season 2019-2020, and their individual models also provide good performance as shown in Table III.
@@ VI Discussion
In this study, we have analyzed various ensemble methods applied to the challenging task of medical data classification. Our primary objective was to develop a robust classifier for the season of 2019-2020, leveraging historical data spanning from 2011 to 2018. The results obtained from our experiments reveal a temporal trend in the performance of individual models. Contrary to the straightforward expectation that more recent data would inherently yield better results, the final weights of the ensemble model exhibited a complex relationship with the season of training data. Notably, the maximum weight was achieved with data from season 2017-2018, highlighting the importance of allowing the model to autonomously discern relevant patterns and relationships. This finding emphasizes the importance of adopting dynamic and data-driven approaches in medical data classification, as the temporal relevance of information may not always align with conventional intuitions.
@@
The evaluation of ensemble methods against baseline models provided valuable insights. Our proposed PVAW method outperformed other techniques. This superior performance can be attributed to the method’s inherent adaptability, which allows it to continuously learn and assign weights to individual models based on their performance on the most recent data. By doing so, our approach effectively captures emerging patterns and evolving relationships within the data, resulting in enhanced classification accuracy. In an environment where medical knowledge and practice are continually evolving, our research underscores the potential of dynamic ensemble methods to provide more accurate and efficient solutions for medical data classification tasks. These findings have significant implications for healthcare decision-making, as they demonstrate the importance of data-driven adaptability in capturing the intricate nuances of medical data.
@@
In spite of the promising results presented in this study, there remain several noteworthy limitations that warrant attention in future research. In our study, while the Naive Bayes models proved effective, they represent only a portion of the potential modeling techniques available for medical data classification. These models, although robust, may not fully capture the complexities of medical data as efficiently as more advanced algorithms, such as neural networks. Therefore, future research should explore incorporating these advanced machine learning techniques, which may offer an improvement in the modeling of the nuanced patterns in medical data. Additionally, the dynamic penalty factor used in our study, though effective, has room for refinement. For example, enhancing its adaptability to the changing data distributions in the target domain could lead to more accurate and responsive models.
@@
Another limitation of our study is the specificity of our dataset, which was centered around RSV information from UPMC. This presents a potential constraint in terms of the generalizability of our findings across different medical contexts. The healthcare field is diverse, and patterns observed in one dataset may not be universally applicable. Our future research will aim to validate the model across a broader spectrum of biomedical scenarios, for example, we can incorporate data from different medical institutions. This would help assess the transferability of our approach and determine necessary modifications for different medical settings. Expanding the scope of our research through collaborative and cross-disciplinary studies could also provide new insights and data sources, enhancing the model’s applicability and robustness.

### 07 Apr 2024 v2

@@ IV Dataset
The research data were sourced from the University of Pittsburgh’s Neptune Research Data Warehouse [43], and the study protocol received approval from the University of Pittsburgh Institutional Review Board (Protocol STUDY20030193). The dataset consists of 2,263,360 encounters from five UPMC EDs (Presbyterian University Hospital, Children’s Hospital, McKeesport Hospital, Mercy Hospital, and Shadyside Hospital) in Allegheny County over nine years from January 2011 to May 2020. In regions with temperate climates, RSV infections predominantly manifest during the colder winter months. The onset of these infections is usually observed around October and November in the Northern Hemisphere. Typically, the highest incidence of RSV cases, or the peak period, occurs in January and February. As the season progresses, the frequency of RSV infections starts to decline, with the final cases of the season generally being recorded between March and May [12]. Given the seasonal nature of RSV outbreaks, we defined one season as spanning from June of one year to May of the following year, rather than using a calendar year division for data segmentation. Table I details the sizes of encounters for each season. An encounter was classified as related to RSV if it included one of any of the following International Classification of Diseases (ICD) codes at discharge; encounters that did not include any of these ICD codes were categorized as non-RSV.
@@
Our dataset contains 7,614 features, including demographic information, lab test results, and NLP-extracted clinical findings, detailed in Table II. The clinical findings were obtained by using the MetaMap parser [44], which applies pattern-matching and deduction rules. The clinical notes for each encounter serve as input for the parser, which extracts a set of clinical findings and their values. The values for each finding are either present (P), negated (N), or missing (M), with missing denoting that clinicians did not mention the finding in the note, as per the parser’s output. The recommendation to designate missing values in NLP-retrieved findings from EHRs as a distinct category was made by [45]. In cases where an encounter had multiple clinical notes, we used the union of all clinical findings across all the notes. Contradictions across notes were resolved in favor of present findings, and this also applied to multiple occurrences of a finding in a single note.
@@ V-A Experiment Settings
We employed an online multi-source transfer learning ensemble framework to enhance model’s adaptability in a specific target domain. This framework incorporated datasets from previous seasons, each treated as a distinct source domain. We evaluated models’ performance using the Area Under the Receiver Operating Characteristic Curve (AUROC), an important measure for gauging a model’s discriminative performance when classifying positive and negative cases.
@@
In the specified algorithm, two hyperparameters, $\alpha$ and $\beta$, play crucial roles in the dynamic adjustment of model weights during the learning process. The setting of $\alpha$ and $\beta$ is consistent across all domains. The hyperparameter $\alpha$ is set to $\log_{2}10$ = 3.32, a value selected after experimenting with several hyperparameter settings. This value for $\alpha$ was chosen to balance the weight penalty applied to a classifier’s weights upon making an incorrect classification. This particular value for $\alpha$ strikes a balance between being too lenient and too punitive, allowing for a gradual decrease in a classifier’s influence while still retaining its contributions where they prove useful.
@@
The hyperparameter $\beta$, set at $1/200000$, plays a critical role in the volume-adaptive component of our algorithm. This parameter is instrumental in fine-tuning the adjustment of the target weight $w_{T}$, aligning it with the growing volume of incoming data. Here, the factor $j$ serves as an index representing the sequence of data points being processed. The specific choice of $1/200000$ for $\beta$ strikes a delicate balance. When the number of instances has not yet reached 200,000, this configuration ensures that the target weights are subtly suppressed, preventing premature dominance in the model’s decision-making process. However, once the count surpasses this threshold, the target weights are progressively amplified, reflecting the increased reliability and relevance of the more substantial data corpus. This ensures that the target model’s influence is appropriately scaled, growing more pronounced as the volume of target data becomes more significant, thereby sensitively responding to the increase in data volume while simultaneously tempering the rate of weight adjustment.
@@
By adopting this value for $\beta$, the algorithm ensures a balanced and responsive adaptation to the accumulating target data. Such a setup allows our ensemble model to systematically transition its focus, increasingly prioritizing the most recent data, which is presumed to be the most relevant for the current learning context. This approach exemplifies the nuanced and progressive nature of online learning, where the model dynamically adapts to the evolving data landscape, ensuring that its classifications remain accurate and pertinent.
@@ V-B Results and Analysis
Our primary objective is to develop a model adept at classifying data from the current season, which is data from the season 2019-2020 in this study. To achieve this, we trained individual models for each season ranging from 2011 to 2018. Subsequently, we evaluated the performance of these models using all of the data from season 2019-2020 to ascertain their effectiveness in handling present-season data. The results of using these hyperparameters are presented in Table III, providing a clear overview of each model’s performance across this temporal span.
@@
For the results, a clear upward trend in AUROC scores is evident as we approach the season of 2019-2020, although 2011-2012 is an outlier. This trend is especially pronounced in models trained with more recent data, indicating an improvement in the classifiers’ ability to distinguish between positive and negative cases as they are trained on data closer to the target season. This improvement can be attributed to the increasing relevance and recency of the data. As the training data gets closer to the target season, it likely becomes more representative of the current patterns and trends, thereby enhancing the classifiers’ accuracy.
@@
To evaluate the performance of MSAW strategy, we compare them with baseline methods, including:
@@
  • Pre-trained Using Source: This approach involves pre-training a model using a combined dataset from all source domains.
  • Online Learning Model: This model employs an online learning strategy, focusing exclusively on sequentially acquired data from the target domain for training. In this approach, the model is tested on an incoming data point and then trained on it before moving on to the next, thus continuously updating its parameters in a test-and-train cycle to adapt to new data.
  • Online Learning Model Pre-trained Using Source: This approach combines two previous methods together by pre-training a model using all source domains and online learning on the target domain.
@@
  • Ensemble Method with Equal Weights: This method employs an ensemble method that integrates various individual source models, assigning them equal and fixed weights throughout the process.
@@
  • Ensemble Method with Volume-based Weights: This method uses an ensemble method that weights individual source models according to the volume of their respective training datasets.
@@
  • Ensemble Method with Time-based Weights: This ensemble method integrates individual models based on their temporal proximity to the testing season (target domain), defining ‘time distance’ as the number of seasons separating a model’s training data from the testing period.
@@
In our comparative analysis of various online learning strategies against the backdrop of sequential epidemiological data, distinct performance metrics emerge, highlighting the efficacy and adaptability of these models as shown in Table IV. The model pre-trained using all of the source domains, at an AUROC of 0.852, demonstrates the potential advantages of a comprehensive, cross-domain dataset for pre-training, indicating the value of extensive historical data. The online learning model, with an AUROC of 0.850, showcases a commendable capacity for leveraging sequentially acquired data, underscoring the viability of online learning in dynamically evolving domains. Online learning started with a pre-trained model boosts the AUROC to 0.864. Conversely, the ensemble with equal weights method registers a lower AUROC of 0.728, suggesting that uniform weighting across diverse models may dilute the classification precision. More notably, the ensemble with volume-based weights, at an AUROC of 0.680, indicates that volume-centric model weighting may not effectively capture the intricate data nuances necessary for accurate target domain classifications. An improvement is observed with the ensemble with time-based weights, achieving an AUROC of 0.761, which implies that temporal relevance plays a crucial role in enhancing model performance, albeit not to the extent of the more sophisticated methods. Our MSAW strategy exhibits an AUROC of 0.870, which is superior to all other methods. This delineates the MSAW strategy’s predominant advantage in dynamically calibrating model weights based on the evolving data landscape, marking a significant stride towards optimizing predictive models for healthcare applications, particularly in the nuanced domain of RSV data analysis.
@@
TABLE IV: Comparative results of different models, featuring P values from DeLong’s test for statistical differences between one AUROC and the AUROC of the MSAW method.
@@
Online Learning Model Pre-trained Using Source 0.864 (P$<$0.001)
@@
MSAW 0.870
Figure 2: (a) Probability of a feature present given RSV
(b) $Log_{10}LR$ of features present.
Our proposed MSAW approach, which utilizes dynamically updated weights, surpasses these conventional baseline methods. The essence of its enhanced performance lies in its ability to dynamically adapt, customizing its relevance to the current data. This dynamic adaptation is achieved by continuously adjusting the weights of the models based on their immediate performance with new data. Such a strategy enables the model to respond effectively to changing data trends and to identify patterns that were not apparent in earlier training datasets. As a result, our method shows an improvement in both accuracy and effectiveness in classifying medical data, outperforming the other methodologies evaluated in the study.
Fig. 2(a) presents a comparison of the probabilities of features present in RSV cases across different seasons, as calculated in our training dataset. Fig. 2(b) compares the $Log_{10}$ likelihood ratios (LR) between RSV cases and non-RSV encounters. A positive LR value suggests that the presence of the feature is more probable in RSV cases than in non-RSV encounters. Among the features for which the probability P(feature||||RSV) exceeded 0.01 in the 2019-2020 season, we identified the top 13 features with the LR. In Fig. 2(a), the probability of ‘Age-less than 5’ group appears high with values ranging from 0.68 to 0.96 in all years, while other features only have around or below 0.2. As shown in Fig. 2(b), with relatively small ranges of values, ‘Age-less than 5’ and ‘Increased work of breathing’ are two significantly steadier features than others observed through multiple seasons. The values in models of seasons 2012-2013 and 2013-2014 are the most divergent from the model of 2019-2020 season, suggesting that the feature patterns in the 2012-2013 and 2013-2014 seasons are the least similar to our target season in 2019-2020, which also aligns with our model results in Table III where seasons 2012-2013 and 2013-2014 present the lowest AUROC. Conversely, the values for models corresponding to the 2017-2018 and 2018-2019 seasons are predominantly found nearest to the 2019-2020 season, and their respective models also demonstrate strong performance, as detailed in Table III.
@@ VI Discussion
In this study, we have analyzed various ensemble methods applied to the challenging task of medical data classification. The primary goal was to develop a robust classifier for the season of 2019-2020, leveraging historical data spanning from 2011 to 2018. The results obtained from our experiments do not necessarily reveal a temporal trend in the performance of individual models. The model developed using data from season 2011-2012 performed much better than season 2012-2013, 2013-2014, and 2014-2015, highlighting the importance of allowing the model to autonomously discern relevant patterns and relationships. This finding emphasizes the importance of adopting dynamic and data-driven approaches in medical data classification, as the temporal relevance of information may not always align with conventional intuitions.
@@
The evaluation of ensemble methods against baseline models provided valuable insights. Although the AUROCs of the ‘Online Learning Model Pre-trained Using Source’ method and our proposed MSAW are comparable, the MSAW is performing slightly better and is more adaptable to situations in which recent data experience a major domain shift. The MSAW might be expected to perform significantly better when there is a large domain shift in the recent past, for example, new strains of RSV that are symptomatically quite different than previous strains. The MSAW method outperformed other techniques. This superior performance can be attributed to the method’s inherent adaptability, which allows it to continuously learn and assign weights to individual models based on their performance on the most recent data. By doing so, our approach effectively captures emerging patterns and evolving relationships within the data, resulting in enhanced classification accuracy. In an environment where medical knowledge and practice are continually evolving, our research underscores the potential of dynamic ensemble methods to provide more accurate and efficient solutions for medical classification tasks. These findings have significant implications for healthcare decision-making, as they demonstrate the importance of data-driven adaptability in capturing the intricate nuances of medical data.
As shown in Fig. 2, many laboratory tests are correlated with RSV presence. For instance, there is a notable correlation between higher platelet counts and the presence of RSV. This correlation is also mentioned in [46], which has shown that children diagnosed with RSV tend to have significantly higher platelet counts compared to those with negative respiratory virus panels or infections caused by other respiratory viruses, including HMPV, influenza A, influenza B, Adenovirus, and Parainfluenza. As another example, our models indicate that patients with RSV often exhibit higher Monocytes counts per 100 leukocytes. One research [47] aligns with our observation, showing that RSV may trigger monocyte activation. Interestingly, additional findings reveal that even normal levels of oxygen saturation and C-reactive protein are associated with RSV. The act of performing these tests, as opposed to not performing these tests, shows a relationship with RSV presence. All these findings highlight the effectiveness of using observational EHRs to detect correlations between lab results and RSV. Importantly, Fig. 2 demonstrates annual variations in the correlation strength between laboratory results and RSV, possibly due in part to adjustments in laboratory testing protocols for different age groups and patient severity levels.
@@
In spite of the promising results presented in this study, there remain several noteworthy limitations that warrant attention in future research. In our study, while the Naive Bayes models proved effective, they represent only one simple modeling technique for medical data classification. For more intricate classification tasks, the Naive Bayes method may not effectively capture the complexities of medical data as well as more advanced algorithms, such as neural networks. Therefore, future research should explore incorporating these advanced machine learning techniques, which may offer an improvement in the modeling of the nuanced patterns in medical data. Additionally, the dynamic penalty factor used in our study, though effective, has room for refinement. For example, enhancing its adaptability to the changing data distributions in the target domain could lead to more accurate and responsive models. Specifically, rather than uniformly penalizing all weights, we aim to explore more sophisticated methods to adjust the weights of the model’s parameters for different features to varying extents. This approach would enable a more accurate reflection of changes.
@@
Another limitation of our study stems from our dataset. RSV cases are identified using ICD codes rather than confirmed through RSV testing results. Consequently, non-RSV encounters might include some RSV-positive cases lacking RSV-related ICD codes. However, given the vast number of non-RSV encounters, this issue is unlikely to significantly impact the results. In addition, the specificity of our dataset, which was centered around RSV information from UPMC, presents a potential constraint in terms of the generalizability of our findings across different medical contexts. The healthcare field is diverse, and patterns observed in one dataset may not be universally applicable. Our future research will aim to validate the model across a broader spectrum of biomedical scenarios. For example, we can incorporate data from different medical institutions, which would help assess the transferability of our approach and determine necessary modifications for different settings. Expanding the scope of our research through collaborative and cross-disciplinary studies could also provide insights and data sources, enhancing the model’s applicability and robustness.